OpenStudy (anonymous):

What is a perfect square divisible by 275? (Simplifying Radicals Homework.)

OpenStudy (amistre64):

275^2 maybe

OpenStudy (cwrw238):

275 = 5*5*11 perfect square = 5*5*11*11

OpenStudy (cwrw238):

= 3025 - thats the lowest perfect square divisible by 275